OpenStudy (macabbey101):

3^2x3^5 I TOTALLY FORGOT HOW TO DO THIS :/

OpenStudy (fibonaccichick666):

what is 3^2?

OpenStudy (squirrels):

\[3^2\] that is 3^2

OpenStudy (anonymous):

(x^a)(x^b) = x^(a+b)

OpenStudy (macabbey101):

thats really confusing lol but i think i can get it

OpenStudy (squirrels):

\[a^b\] "b" means to multiply the same number (a) that many times.

OpenStudy (macabbey101):

yeah, i know that part, but i forgot how to multiple two like that :/ get what im saying?

OpenStudy (fibonaccichick666):

whoops sorry, if you have the same base and you are multiplying, do you add or multiply the exponents to simplify the expression?

OpenStudy (squirrels):

for example \[4^3\] = 4*4*4

OpenStudy (fibonaccichick666):

Yea, I know sorry dislexic at times

OpenStudy (squirrels):

You get what im saying?

OpenStudy (fibonaccichick666):

@Squirrels that's not the question

OpenStudy (squirrels):

i know, im giving him an example

OpenStudy (macabbey101):

im a girl lol

jimthompson5910 (jim_thompson5910):

When multiplying exponential expressions with the same base, you add the exponents. Example: \[\Large 4^3 \times 4^5 = 4^{3+5} = 4^{9}\]

OpenStudy (macabbey101):

thats okay though, i get what you are saying haha :) @Squirrels

OpenStudy (macabbey101):

THANK YOU!!:) @jim_thompson5910

OpenStudy (squirrels):

ok good haha

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (squirrels):

sorry i mean her* lol

OpenStudy (macabbey101):

thats okay :)

jimthompson5910 (jim_thompson5910):

oops typo, meant to say \[\Large 4^3 \times 4^5 = 4^{3+5} = 4^{8}\] but you get the idea